OpenStudy (anonymous):

what is the domain of 4(x^2-100)

OpenStudy (luigi0210):

have you tried it?

OpenStudy (dumbcow):

domain is everything except the vertical asymptotes....so basically this question is find the vertical asymptotes

OpenStudy (anonymous):

so i would do 4x^2-400=0?

OpenStudy (anonymous):

would it be -5 & 5?

OpenStudy (anonymous):

or would is be 10 & -10?

OpenStudy (dumbcow):

crap i go too fast ... is this a rational function \[y = \frac{4}{x^{2} -100}\] ??

OpenStudy (dumbcow):

btw it would be +-10

OpenStudy (dumbcow):

hello... is it a fraction or is it \[y = 4(x^{2} -100)\] ? in which case domain is all real num

OpenStudy (anonymous):

it was 10 ,-10 :)

OpenStudy (dumbcow):

:)